OpenStudy (anonymous):

2x + 15 - 3x when x = 4?

OpenStudy (anonymous):

2 * 4 = 8 3 * 4 = 12 8 + 15 - 12 = what?

OpenStudy (anonymous):

11

OpenStudy (anonymous):

There you go(:

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

No problem :)